Suppose that two conformations A and B of a molecule differ in energy by 5.0 kJ/ mol, and a third conformation lies 0.50 kJ/mol above B. What proportion of molecules will be in conformation B at 273 K, with each conformation treated as a single energy level?
Answer #1

Answer:

Let NA = molecules in A level

Let NB = molecules in B level

Let NC = molecules in C level

According to Boltzmann distribution

NB/NA = exp(-(EB-EA)/RT) = exp(-5.0 x 1000J / 8.314 x 273) = 0.111

NC/NB = exp(-(EC-EB)/RT) = exp(-0.5 x 1000J / 8.314 x 273) = 0.803

Also NA + NB + NC = NA + 0.111 NA + NC = NA + 0.111 NA + 0.803 NB =   NA + 0.111 NA + 0.803 x 0.111 NA = 1.200 NA = No

So No = 1.200 NA     so NA = (1.0/1.200) No

NB = 0.111 NA = 0.111 x (1.0/1.200) No = 0.0925 No

So percentage abundance of B state is = (NB /No) x 100 = 0.0925 x100 = 9.25 %